The method of substitution to solve the system of linear equations.
3x-y=3
7x-2y=9
I think i am suppose to first get y by it self
y=3x-3
then I am suppose to divide 3 from both sides?
y=3x/3-3/3
I think this is it..
y=3x-3
7x-2y=9
get y on the other side
2y=7x-9
am I suppose to subtract 2 from both ends?
y=7x-?
I am lost form here.

Answer by Alan3354(69443)

The method of substitution to solve the system of linear equations.
3x-y=3
7x-2y=9
I think i am suppose to first get y by it self
y=3x-3
then I am suppose to divide 3 from both sides?
y=3x/3-3/3
I think this is it..
y=3x-3
================
That's right, solve for y (in this problem. Sometimes you solve for x)
You don't divide by 3, tho, just solve for y.
y=3x-3
-------
Then sub for y in the other equation.
----
7x-2y=9
7x - 2(3x-3) = 9
Solve for x
7x -6x+6 = 9
x = 3
========
y = 3x-3
y = 6
That's all.
=====================
=====================
get y on the other side
2y=7x-9
am I suppose to subtract 2 from both ends?
y=7x-?
I am lost form here.